OpenStudy (anonymous):

hi can you please help me factored form for n^2+16n+64

hero (hero):

Find two numbers \(p\) and \(q\) such that \(p \times q = 64\) \(p + q = 16\)

OpenStudy (vjordillo):

It is an example of a perfect square trinomial. So, the factors are (x+8)(x+8)

OpenStudy (anonymous):

what about this one n^2+24n+144

OpenStudy (anonymous):

can you explain too please

OpenStudy (vjordillo):

*this is true to perfect square trinomials only* First, let y= be a number (x+y)(x+y)=0 Then, expand the given terms, \[x ^{2}+2yx+y^2=0\] Since for a quadratic function \[f(x)=Ax ^{2}+Bx+C\] Your A=1, B=2y and C=y^2

OpenStudy (vjordillo):

Your y^2 in your equation is 144, therefore the value of your y=12. Substituting 12 to the as your b will make 24 as the coefficient of x. The factors will just be \[(x+12)^{2}\]

OpenStudy (vjordillo):

*substituting 12 as your b ...